OpenStudy (anonymous):

-5/4x = 15/16. It is -3/4 or 77/64

OpenStudy (a_clan):

x is in numerator or denominator?

OpenStudy (anonymous):

it's more like -4/3.... the way it's given

OpenStudy (a_clan):

if it's in numerator, -3/4 if in denominator , -4/3

OpenStudy (anonymous):

-5/4X=15/16 DIVIDE BOTH SIDES BY -5 -5/4x*1/-5=15/16*1/-5 1/4x=-3/16 multiply both sides by 4 (1/4x)*4=(3/16)*4 1/x=3/4 cross multiplication 3x=4 x=4/3

OpenStudy (anonymous):

kushashwa... you lost "-" :)

OpenStudy (annon):

x=-300/16

OpenStudy (anonymous):

why inik

OpenStudy (annon):

sorry it's -3/4